# Angular-momentum pairs in spherical systems: applications to the Galactic centre

### v1

@@ 3.1 Synchronous circular orbits
As we show in Appendix C, the torque between the pair (for small relative angles $\theta_{b}\ll 1$), is given by
$\Omega_{i,j}=2\pi\omega_{i}\frac{m_{j}}{M}\frac{1}{\pi\theta_{b}^{2}}.$ (20)
[…]
@@
$\frac{r}{r_{\bullet}}\geq\begin{cases}\displaystyle\left(\frac{4}{3\pi\kappa}\right)^{1/3}\left(\frac{2m_{i}m_{j}}{m_{\bullet}(m_{i}+m_{j})}\right)^{1/3}\theta_{b}^{-2/3}\lvert\sin\theta_{\bullet}\rvert^{-2/3},&r<r_{\bullet},\\[11.99998pt] \displaystyle\left(\frac{3\pi\kappa}{4}\right)^{1/2}\left(\frac{2m_{i}m_{j}}{m_{\bullet}(m_{i}+m_{j})}\right)^{-1/2}\theta_{b}\lvert\sin\theta_{\bullet}\rvert,&r\geq r_{\bullet}.\end{cases}$ (22)
@@
Figure 2 confirms the validity of condition (19). In the top rows of both panels (a) and (b), the blue curves show the torques within each pair (the left-hand side of Eq. (19) multiplied by $|\bm{L_{i}}|/m_{i}$), while the red curves show the differential torques exerted by the perturber (the right-hand side of Eq. (19) multiplied by $|\bm{L_{i}}|/m_{i}$). Each column corresponds to an initial mutual inclination between the two objects of $4^{\circ}$, $7^{\circ}$, and $10^{\circ}$, respectively for $\kappa=0.5$. According to condition (19), a pair becomes unbound where the red curve lies above the blue one. The bottom rows of both panels present N-Ring simulations for the same initial configurations. Each pair was placed uniformly on the unit sphere with initial mutual inclinations of $4^{\circ}$, $7^{\circ}$, and $10^{\circ}$, and the three rings were evolved according to the equations of motion (7). The panels display the mutual inclination of each pair versus its initial inclination to the perturber at four epochs: $t=0$, 1, 5, and $10\,\mathrm{Myr}$. The numerical results align beautifully with the analytic prediction: all pairs remain bound in the $4^{\circ}$ case (left column), whereas those located above the blue–red intersection in the $7^{\circ}$ and $10^{\circ}$ cases become unbound.
@@
Fig. 2(a) illustrates the quadrupole-only prediction obtained by retaining solely the $\ell=2$ term and adopting a radius ratio $r/r_{\bullet}=a/a_{\bullet}=1/3$. We clearly see one peak for the differential torque at $\theta_{\bullet}=90^{\circ}$. By contrast, Fig. 2(b) displays the outcome when the multipole expansion is summed to $\ell_{\max}=200$ at the larger separation $r/r_{\bullet}=a/a_{\bullet}=2/3$. The additional harmonics introduce the secondary peaks visible in the red perturber-curve, yet in both panels the crossing of the blue (pair) and red (perturber) curves still marks the critical boundary, beyond which the N-Ring simulations in the lower rows show the binaries becoming unbound.
@@
Additionally, Fig. 2(a) illustrates a case with relatively fast nodal precession $\Omega_{\bullet}^{-1}(\pi/4,a/a_{\bullet})\approx 3.5\times 10^{5}$ yr about the angular momentum of the perturber with mass $m_{\bullet}=5\times 10^{4}\,M_{\odot}$. The figure shows that for the initial pair’s inclination of 7, the disrupted pairs reached maximum separation at $t=5$ Myr (orange curve in the second row), while at 10 Myr (green curve) the angular separations are low again, illustrating the oscillatory behaviour for the disrupted pairs between $\theta_{b}$ and $\theta_{1}+\theta_{2}$, as evident from the geometry of the problem (see Fig. 1). A similar pattern is observed for the initial angular separation of $10^{\circ}$, but with a slightly faster disruption time-scale. On the other hand, Fig. 2(b) depicts the case where nodal precession is much slower, $\Omega_{\bullet}^{-1}(\pi/4,a/a_{\bullet})\simeq 20$ Myr for a perturber with mass $m_{\bullet}=2\times 10^{3}\,M_{\odot}$. The orange and green curves show that pairs with low-inclination orbits have already reached their maximum separation, whereas those that start at higher inclinations are still progressing toward that point.
@@ 3.2 Hierarchical mass-ratio pairs
Let us consider the case where $a_{i}=a_{j}$ as previously, but with $m_{i}\gg m_{j}$. In this case, $\bm{L}_{j}$ precesses about $\bm{L}_{i}+\bm{L}_{j}\approx\bm{L}_{i}$ in the absence of an IMBH, while conserving $\hat{\bm{L}}_{i}\cdot\hat{\bm{L}}_{j}$; this is thus a nodal precession about $\bm{L}_{i}$. The analogy with tides (equivalently, the restricted three-body problem) implies that equation (15) should reduce to a comparison between the torque $\bm{\Omega}_{ij}\times\bm{L}_{j}$ and the torque-difference $\Omega_{j,\bullet}\times\bm{L}_{j}$ between opposite positions in the orbit of $\bm{L}_{j}$ about $\bm{L}_{i}$ – when the argument of the ascending node of particle $j$ is 00 and $\pi$. We denote the difference between the corresponding values of $\bm{L}_{j}$ by $\delta\bm{L}_{j}$, which is necessarily perpendicular to $\bm{L}_{i}$. The above reasoning still applies, and therefore the stability condition for this pair is still
@@
3.3 Elliptical orbits
Figure 4: Contour lines show the analytical boundary of stability for a pair of circular orbits against disruption by an eccentric IMBH for initial mutual inclination of $7^{\circ}$, i.e. regions where the torque between a pair of bodies equals the right hand side of the Eq. (19) for eccentric orbits, calculated from the analytical expressions in section 3.3. Vertical lines show the pericentre, the semi-major axis, and the apocentre of the IMBH. The IMBH mass, eccentricity, and $\kappa$ parameter are indicated in the panels.
In general the orbits can be divided into two basic geometrical classes: radially non-overlapping ($r_{\mathrm{a}}<r_{\rm p,\bullet}$ or $r_{\mathrm{p}}>r_{\rm a,\bullet}$) and radially overlapping orbits ($r_{\rm a}\geq r_{\rm p,\bullet}\text{~{}and~{}}r_{\rm p}\leq r_{\rm a,\bullet}$) with respect to the perturber. The mutual torques are strongest for overlapping orbits (see Kocsis & Tremaine 2015), whence, as for the circular case, we always consider only overlapping orbits for the pair, and use the asymptotic expressions to describe mutual torque within the pair. Thus, the angular frequency can be approximated as:
@@ 3.3.1 Non-overlapping orbits with respect to the perturber
which can be simplified when both components of the pair are on circular orbits with the same radii:
$\theta_{b}\;\geq\;\begin{cases}\displaystyle\left(\frac{4}{3\pi\kappa}\right)^{1/2}\left(\frac{2m_{i}m_{j}}{m_{\bullet}(m_{i}+m_{j})}\right)^{1/2}\left(\frac{r}{a_{\bullet}}\right)^{-3/2}\frac{1}{(1-e^{2}_{\bullet})}\frac{1}{\lvert\sin\theta_{\bullet}\rvert}.&r<a_{\bullet}\\[13.99995pt] \displaystyle\left(\frac{4}{3\pi\kappa}\right)^{1/2}\left(\frac{2m_{i}m_{j}}{m_{\bullet}(m_{i}+m_{j})}\right)^{1/2}\left(\frac{a_{\bullet}}{r}\right)\left(\frac{1-e^{2}_{\bullet}}{P_{3}(\chi_{\bullet})}\right)^{1/2}\frac{1}{\lvert\sin\theta_{\bullet}\rvert}.&r\geq a_{\bullet}\end{cases}$ (27)
@@ 3.3.2 Overlapping orbits with the perturber
$\theta_{b}>\frac{1}{\kappa}\frac{2m_{i}m_{j}}{m_{\bullet}(m_{i}+m_{j})}\frac{a_{\bullet}}{a_{i}}\frac{I_{2}^{(i,j)}}{I_{2}^{(i,\bullet)}}|\sin{\theta_{\bullet}}|$ (28)
Note that due to complicated dependence of $I_{2}$ on the pericentre and apocentre distances, the expression can be explored only numerically to solve for the semi-major-axis ratio with respect to the IMBH. The dependence on the eccentricities of the pair and of the perturber is also contained implicitly in $I_{2}$. Note that $I_{2}/a_{\rm out}$ is shown in Kocsis & Tremaine (2015), bottom panels of Figure 1 therein, which exhibit two narrow singular spikes as a function of semi-major axis and eccentricity, when the periapsis and apoapsis of the pair and the perturber respectively coincide, and has a local minimum in between these spikes, and is monotonic in the region outside of both spikes.
@@
$\theta_{b}>\frac{1}{\sqrt{\kappa}}\frac{\pi}{2}\left(\frac{2m_{i}m_{j}}{m_{\bullet}(m_{i}+m_{j})}\right)^{1/2}\frac{a_{\bullet}^{1/2}\left[(r-r_{\mathrm{p,\bullet}})(r_{\mathrm{a,\bullet}}-r)\right]^{1/4}}{r}\sqrt{|\sin{\theta_{\bullet}}|}.$ (29)
@@
Figure 5: Average inclination angle between pairs after $5$ Myr (two top rows) and $10$ Myr (two bottom rows) of evolution as a function of relative distance from the IMBH and relative inclination with respect to the IMBH. The colour coding represents the average mutual inclination. Odd rows – N-Ring, even rows – $\phi$-CPU, for the identical parameters.
@@
The $N$-body integrations (rows 2 and 4) exhibit disruption over a broader extent of the $(\theta_{\bullet},a)$ plane and on shorter time-scales than the N-Ring models, with irregular – but recognisable – boundaries that echo the analytic contours in Fig. 4. The bright “islands” near the IMBH pericentre, semi-major axis, and apocentre arise from higher multipole terms, as predicted. Notably, in the hierarchical mass-ratio configuration (column iii) the N-Ring runs show no disruption, consistent with the secular criterion, whereas $\phi$-CPU reveals a sizeable unstable region. These discrepancies stem from resonant effects and energy–angular-momentum exchange – both absent in the orbit-averaged formalism—which alter the semi-major axes and eccentricities in the full $N$-body dynamics, and are prominent here due to the large mass of the pair. We quantify the regime of validity of the VRR approximation in the next section, but emphasise that the overall morphology of the disruption regions remains in good qualitative agreement across the analytic, N-Ring, and $N$-body methods, supporting the robustness of our model.
@@ 4.1 Three-body 1:1 resonance
The level set of $H_{r}$ are plotted in the right panel of Figure 6. The resonance is at $(\varphi,J_{r})=(0,0)$; there, $\cos\varphi=1$, so the potential term diverges at first order in $J_{r}$. This requires an expansion to second order in $J_{r}$ in the denominator, which shows that in fact this term scales like $\varepsilon/J_{r}$ near the resonance. The width of the resonant region is obtained by studying the level set of $H_{r}$: libration occurs when the star is effectively bound to the IMBH – when $\varphi$ is bounded, i.e. for $H_{r}\leq-\varepsilon/2$ – and circulation occurs when $H_{r}\geq-\varepsilon/2$; the line $H_{r}=-\varepsilon/2$ is thus the separatrix. The maximum deviation of $\left|J_{r}\right|$ from 00 inside the resonant region is at $\varphi=0$, given by $\left(\varepsilon/3\right)^{1/3}$ – this is just obtained by setting $H_{r}=-\varepsilon/2$ and $\varphi=0$ in equation (4.1) and solving for $J_{r}$ – so that the width of the resonance is
@@
Just outside the separatrix, $J_{r}$ reaches its maximum at $\varphi=0$, and its minimum at ±$\pm\pi$. These are given by
@@
Observe, that if we set $\left|1-\alpha^{2}\right|=(\varepsilon/3)^{1/3}$ in equation (37), we find $\Delta J_{r}=(\varepsilon/3)^{1/3}$ – exactly half of the librating variation. The factor of two is expected, $J_{r}$ does not cross 00 outside the separatrix.
Figure 6: Left: relative changes of Keplerian energies for all pairs from $N$-body simulations over 10 Myr. Physical units are the same as in Fig. 2(b). The systematic increase in energy with distance ratio is due to the assumed Plummer potential which represents the contribution of stars. Right: a level set of the Hamiltonian $H_{r}$ (equation (4.1)), as a function of the resonant action $J_{r}=a^{2}/a_{\bullet}^{2}-1$ and the angle $\varphi$ (for $\varepsilon=m_{\bullet}/M=0.1$). The ordinate is re-scaled by $\varepsilon^{1/3}$, as found in equation (34).

### v2

@@ 3.1 Synchronous circular orbits
As we show in Appendix C, the angular frequency between the pair (for small relative angles $\theta_{b}\ll 1$), is given by
$\Omega_{i,j}=2\omega_{i}\frac{m_{j}}{M}\frac{1}{\theta_{b}^{2}}.$ (21)
Substituting the Eq. (21) for the left-hand-side of the Eq. (19) the quadrupole approximation in Eq. (2.2) for the right-hand into the disruption condition (19) and retaining the small-angle approximation $\sin\theta_{b}\simeq\theta_{b}$, we determine the conditions under which the pair decouples due to the perturber’s influence. As remarked above, the inequality (19) can be re-arranged to solve for the relative angle $\theta_{b}$ and the distance ratio $\bm{r}/r_{\bullet}$. First of all, solving for the relative angle $\theta_{b}$, we obtain:
@@
$\begin{cases}\displaystyle\frac{r}{r_{\bullet}}\geq\left(\frac{4}{3\pi\kappa}\right)^{1/3}\left(\frac{2m_{i}m_{j}}{m_{\bullet}(m_{i}+m_{j})}\right)^{1/3}\theta_{b}^{-2/3}\lvert\sin\theta_{\bullet}\rvert^{-2/3},&r<r_{\bullet},\\[11.99998pt] \displaystyle\frac{r}{r_{\bullet}}\leq\left(\frac{3\pi\kappa}{4}\right)^{1/2}\left(\frac{2m_{i}m_{j}}{m_{\bullet}(m_{i}+m_{j})}\right)^{-1/2}\theta_{b}\lvert\sin\theta_{\bullet}\rvert,&r\geq r_{\bullet}.\end{cases}$ (23)
@@
Figure 2 confirms the validity of condition (19). In the top rows of both panels (a) and (b), the blue curves show the torques within each pair (the left‑hand side of Eq. (19) multiplied by $|\bm{L_{i}}|/m_{i}$), while the red curves show the differential torques exerted by the perturber (the right‑hand side of Eq. (19) multiplied by $|\bm{L_{i}}|/m_{i}$). Each column corresponds to an initial mutual inclination between the two objects of $4^{\circ}$, $7^{\circ}$, and $10^{\circ}$, respectively for $\kappa=0.5$. According to condition (19), a pair becomes unbound where the red curve lies above the blue one. The bottom rows of both panels present N-Ring simulations for the same initial configurations. Each pair was placed uniformly on the unit sphere with initial mutual inclinations of $4^{\circ}$, $7^{\circ}$, and $10^{\circ}$, and the three rings were evolved according to the equations of motion (7). The panels display the mutual inclination of each pair versus its initial inclination to the perturber at four epochs: $t=0$, 1, 5, and $10\,\mathrm{Myr}$. The numerical results align beautifully with the analytic prediction: all pairs remain bound in the $4^{\circ}$ case (left column) [^2] We note that here, the analytic curve for the differential torque marginally exceeds the internal torque near $\theta_{\bullet}=90^{\circ}$, which would formally imply instability. The fact that the simulations show these pairs remaining bound indicates that the effective geometric factor is slightly smaller than the adopted value of $\kappa=0.5$., whereas those located above the blue–red intersection in the $7^{\circ}$ and $10^{\circ}$ cases become unbound.
@@
Fig. 2(a) illustrates the quadrupole‐only prediction obtained by retaining solely the $\ell=2$ term and adopting a radius ratio $r/r_{\bullet}=a/a_{\bullet}=1/3$. We clearly see one peak for the differential torque at $\theta_{\bullet}=90^{\circ}$. By contrast, Fig. 2(b) displays the outcome when the multipole expansion is summed to $\ell_{\max}=200$ at the larger separation $r/r_{\bullet}=a/a_{\bullet}=2/3$. The additional harmonics introduce the secondary peaks visible in the red perturber-curve, yet in both panels the crossing of the blue (pair) and red (perturber) curves still marks the critical boundary, beyond which the N-Ring simulations in the lower rows show the binaries becoming unbound.
@@
Additionally, Fig. 2(a) illustrates a case with relatively fast nodal precession $\Omega_{\bullet}^{-1}(\pi/4,a/a_{\bullet})\approx 3.5\times 10^{5}$ yr about the angular momentum of the perturber with mass $m_{\bullet}=5\times 10^{4}\,M_{\odot}$. The figure shows that for the initial pair’s inclination of 7, the disrupted pairs reached maximum separation at $t=5$ Myr (orange curve in the second row), while at 10 Myr (green curve) the angular separations are low again, illustrating the oscillatory behaviour for the disrupted pairs between $\theta_{b}$ and $\theta_{1}+\theta_{2}$ ($2\pi-(\theta_{1}+\theta_{2})$ when $\theta_{1}+\theta_{2}>180^{\circ}$), as evident from the geometry of the problem (see Fig. 1). A similar pattern is observed for the initial angular separation of $10^{\circ}$, but with a slightly faster disruption time-scale. On the other hand, Fig. 2(b) depicts the case where nodal precession is much slower, $\Omega_{\bullet}^{-1}(\pi/4,a/a_{\bullet})\simeq 20$ Myr for a perturber with mass $m_{\bullet}=2\times 10^{3}\,M_{\odot}$. The orange and green curves show that pairs with low‐inclination orbits have already reached their maximum separation ($\cos\theta_{b}\lesssim\cos(\theta_{1}+\theta_{2})$), whereas those that start at higher inclinations are still progressing toward that point ($\cos\theta_{b}\ll\cos(\theta_{1}+\theta_{2})$), especially the ones closer to the equatorial plane.
@@ 3.2 Hierarchical mass-ratio pairs
Let us consider the case where $a_{i}=a_{j}$ as previously, but with $m_{i}\gg m_{j}$. In this case, $\bm{L}_{j}$ precesses about $\bm{L}_{i}+\bm{L}_{j}\approx\bm{L}_{i}$ in the absence of an IMBH, while conserving $\hat{\bm{L}}_{i}\cdot\hat{\bm{L}}_{j}$; this is thus a nodal precession about $\bm{L}_{i}$. The analogy with tides (equivalently, the restricted three-body problem) implies that equation (15) should reduce to a comparison between the torque $\bm{\Omega}_{ij}\times\bm{L}_{j}$ and the torque-difference $\Omega_{j,\bullet}\times\bm{L}_{j}$ between opposite positions in the orbit of $\bm{L}_{j}$ about $\bm{L}_{i}$ – when the argument of the ascending node of particle $j$ is 0 and $\pi$. We denote the difference between the corresponding values of $\bm{L}_{j}$ by $\delta\bm{L}_{j}$, which is necessarily perpendicular to $\bm{L}_{i}$. The above reasoning still applies, and therefore the stability condition for this pair is still
@@
In Fig. 3 we plot the mean mutual inclination $\langle\theta_{b}\rangle$ in the $(\cos\theta_{\bullet},\;a/a_{\bullet})$ plane at t=5, 10,t=5,\;10, and $20$ Myr (three columns). The top row shows the equal-mass model, the bottom row – the hierarchical model; as expected, the two are nearly indistinguishable. For each panel the dashed contour marks $\langle\theta_{b}\rangle=15^{\circ}$; regions to left of this curve remain effectively bound. The colour scale traces the growth of the pair separation with time and highlights the migration of the differential-torque maxima: moving from larger to smaller $a/a_{\bullet}$ (right to left), the two peaks drift toward $\theta_{\bullet}=0^{\circ}$ and $180^{\circ}$. At the smallest radii the dashed contour is single-peaked, signalling the transition to the quadrupole-dominated regime.
3.3 Elliptical orbits
Figure 4: Contour lines show the analytical boundary of stability for a pair of circular orbits against disruption by an eccentric IMBH for initial mutual inclination of $7^{\circ}$, i.e. regions where the torque between a pair of bodies equals the right hand side of the Eq. (19) for eccentric orbits, calculated from the analytical expressions in section $3.$ Vertical lines show the pericentre, the semi-major axis, and the apocentre of the IMBH. The IMBH mass, eccentricity, and $\kappa$ parameter are indicated in the panels.
@@ 3.3.1 Non-overlapping orbits with respect to the perturber
which can be simplified when both components of the pair are on circular orbits with the same radii: When both components of the pair are on circular orbits with the same radii and eccentric perturber, we also use Eq. (9) for the perturber, but Eq. (88) for the pair to get [^3] Note that the asymptotic expressions used for the torques between pairs behave differently for circular-circular, circular-eccentric and eccentric-eccentric orbital configurations. See Appendix A and Appendix B5 of Kocsis and Tremaine (2015):
$\theta_{b}\;\geq\;\begin{cases}\displaystyle\left(\frac{4}{3\pi\kappa}\right)^{1/2}\left(\frac{2m_{i}m_{j}}{m_{\bullet}(m_{i}+m_{j})}\right)^{1/2}\left(\frac{r}{a_{\bullet}}\right)^{-3/2}\frac{1}{(1-e^{2}_{\bullet})}\frac{1}{\lvert\sin\theta_{\bullet}\rvert},&r<a_{\bullet}\\[13.99995pt] \displaystyle\left(\frac{4}{3\pi\kappa}\right)^{1/2}\left(\frac{2m_{i}m_{j}}{m_{\bullet}(m_{i}+m_{j})}\right)^{1/2}\left(\frac{a_{\bullet}}{r}\right)\frac{(1-e^{2}_{\bullet})^{-1/2}}{\sqrt{P_{3}(\chi_{\bullet})}}\frac{1}{\lvert\sin\theta_{\bullet}\rvert},&r\geq a_{\bullet}.\end{cases}$ (28)
@@ 3.3.2 Overlapping orbits with the perturber
$\theta_{b}\geq\frac{1}{\kappa}\frac{2m_{i}m_{j}}{m_{\bullet}(m_{i}+m_{j})}\frac{a_{\bullet}}{a_{i}}\frac{I_{2}^{(i,j)}}{I_{2}^{(i,\bullet)}}|\sin{\theta_{\bullet}}|$ (29)
Note that due to complicated dependence of $I_{2}$ on the pericentre and apocentre distances, the expression can be explored only numerically to solve for the semi-major-axis ratio with respect to the IMBH. The dependence on the eccentricities of the pair and of the perturber is also contained implicitly in $I_{2}$. Note that $I_{2}/a_{\rm out}$ is shown in Kocsis and Tremaine (2015), bottom panels of Figure 1 therein, which exhibit two narrow singular spikes as a function of semi-major axis and eccentricity, when the periapsis and apoapsis of the pair and the perturber respectively coincide, and has a local minimum in between these spikes, and is monotonic in the region outside of both spikes.
@@
$\theta_{b}\geq\frac{1}{\sqrt{\kappa}}\frac{\pi}{2}\left(\frac{2m_{i}m_{j}}{m_{\bullet}(m_{i}+m_{j})}\right)^{1/2}\frac{a_{\bullet}^{1/2}\left[(r-r_{\mathrm{p,\bullet}})(r_{\mathrm{a,\bullet}}-r)\right]^{1/4}}{r}\sqrt{|\sin{\theta_{\bullet}}|}.$ (30)
@@
Figure 5: Average inclination angle between pairs after $5$ Myr (two top rows) and $10$ Myr (two bottom rows) of evolution as a function of relative distance from the IMBH and relative inclination with respect to the IMBH. In all models we used $a_{\bullet}=0.15$ pc and $m_{\bullet}=2\times 10^{3}\,{\rm M}_{\odot}$. The colour coding represents the average mutual inclination. Odd rows – N-Ring, even rows – $\phi$-CPU, for the identical parameters. The dashed lines show the region where $\theta_{b}<10^{\circ}$. The dotted lines in the third row show the disruption threshold computed using Eq. (19).
@@
The $N$-body integrations (rows 2 and 4) exhibit disruption over a broader extent of the $(\theta_{\bullet},a)$ plane and on shorter time-scales than the N-Ring models, with irregular – but recognisable – boundaries that echo the analytic contours in Fig. 4. The bright “islands” near the IMBH pericentre, semi-major axis, and apocentre arise from higher multipole terms, as predicted. Notably, in the hierarchical mass-ratio configuration (column iii) the N-Ring runs show no disruption, consistent with the secular criterion, whereas $\phi$-CPU reveals a sizeable unstable region. These discrepancies stem from resonant effects and energy–angular-momentum exchange – both absent in the orbit-averaged formalism – which alter the semi-major axes and eccentricities in the full $N$-body dynamics, and are prominent here due to the large mass of the pair. We quantify the regime of validity of the VRR approximation in the next section, but emphasise that the overall morphology of the disruption regions remains in good qualitative agreement across the analytic, N-Ring, and $N$-body methods, supporting the robustness of our model.
@@ 4.1 Three-body 1:1 resonance
The level set of $H_{r}$ are plotted in Figure 6. The resonance is at $(\varphi,J_{r})=(0,0)$; there, $\cos\varphi=1$, so the potential term diverges at first order in $J_{r}$. This requires an expansion to second order in $J_{r}$ in the denominator, which shows that in fact this term scales like $\varepsilon/J_{r}$ near the resonance. The width of the resonant region is obtained by studying the level set of $H_{r}$: libration occurs when the star is effectively bound to the IMBH – when $\varphi$ is bounded, i.e. for $H_{r}\leq-\varepsilon/2$ – and circulation occurs when $H_{r}\geq-\varepsilon/2$; the line $H_{r}=-\varepsilon/2$ is thus the separatrix. The maximum deviation of $\left|J_{r}\right|$ from 0 inside the resonant region is at $\varphi=0$, given by $\left(\varepsilon/3\right)^{1/3}$ – this is just obtained by setting $H_{r}=-\varepsilon/2$ and $\varphi=0$ in equation (4.1) and solving for $J_{r}$ – so that the width of the resonance is
@@
Just outside the separatrix, $J_{r}$ reaches its maximum at $\varphi=0$, and its minimum at $\pm\pi$. These are given by
@@
Observe, that if we set $\left|1-\alpha^{2}\right|=(\varepsilon/3)^{1/3}$ in equation (38), we find $\Delta J_{r}=(\varepsilon/3)^{1/3}$ – exactly half of the librating variation. The factor of two is expected, $J_{r}$ does not cross 0 outside the separatrix.
Figure 6: A level set of the Hamiltonian $H_{r}$ (equation (4.1)), as a function of the resonant action $J_{r}=a^{2}/a_{\bullet}^{2}-1$ and the angle $\varphi$ (for $\varepsilon=m_{\bullet}/M=0.1$). The ordinate is re-scaled by $\varepsilon^{1/3}$, as found in equation (35).
Figure 7: Relative changes of Keplerian energies for all pairs from NN-body simulations over 10 Myr. Physical units are the same as in Fig. 2(b). The systematic increase in energy with distance ratio is due to the assumed Plummer potential which represents the contribution of stars. Left: circular orbits; right: eccentric orbits, with $e_{\bullet}=0.7$ and $e=0.3$ for the stars.
@@
For eccentric orbits, there will be a multitude of resonances, which preclude analytical treatment. The condition found above for there to be a large energy change is to be inside the resonance, delineated by the Hill radius. For the eccentric case, we expect every semi-major axis such that the orbits of the stars and the IMBH overlap, to be inside some resonance, so if $r_{\rm a,\bullet}>r_{\rm p}$ and $r_{\rm p,\bullet}<r_{\rm a}$, then we expect there to be a significant energy exchange, leading to a breakdown of the averaged theory. Then, each value of $a$ that falls between $a_{\bullet}(1+e_{\bullet})/(1-e)$ and $a_{\bullet}(1-e_{\bullet})/(1+e)$, there should be a $\Delta a\sim\varepsilon^{1/3}$. We test that in the right panel of Figure 7, where $e_{\bullet}=0.7$ and $e=0.3$. The condition above implies that every $a\in[3/13,17/7]a_{\bullet}\approx[0.23,2.43]a_{\bullet}$ is strongly affected, which agrees with the simulation data. We also see in this figure, that while for the circular case there exists only a slight imbalance in energy changes as a function of inclination, this is much more pronounced in the eccentric case. Qualitatively, this arises from the known tendency of retrograde orbits to be more stable than prograde ones (Mardling and Aarseth, 2001).